# Towards Real World Debiasing: A Fine-grained Analysis On Spurious Correlation

### v2

@@ Abstract
Spurious correlations in training data significantly hinder the generalization capability of machine learning models when faced with distribution shifts in real-world scenarios. To tackle the problem, numerous debias approaches have been proposed and benchmarked on datasets intentionally designed with severe biases. However, it remains to be asked: 1. Do existing benchmarks really capture biases in the real world? 2. Can existing debias methods handle biases in the real world? To answer the questions, we revisit biased distributions in existing benchmarks and real-world datasets, and propose a fine-grained framework for analyzing dataset bias by disentangling it into the magnitude and prevalence of bias. We observe and theoretically demonstrate that existing benchmarks poorly represent real-world biases. We further introduce two novel biased distributions to bridge this gap, forming a nuanced evaluation framework for real-world debiasing. Building upon these results, we evaluate existing debias methods with our evaluation framework. Results show that existing methods are incapable of handling real-world biases. Through in-depth analysis, we propose a simple yet effective approach that can be easily applied to existing debias methods, named Debias in Destruction (DiD). Empirical results demonstrate the superiority of DiD, improving the performance of existing methods on all types of biases within the proposed evaluation framework.
@@ 1 Introduction
To tackle the problem, various debias methods [13, 12, 14, 9, 15, 16, 17, 18] have been proposed in recent years. And the effectiveness of the methods is benchmarked with synthetic [19, 12, 16] and semi-synthetic [9, 12, 17] (referred as "real-world dataset" in previous works) datasets that are designed to be severely biased. However, while these benchmarks are indeed biased, they are rough and lack thorough consideration of how data is truly biased in the real world. This raises two questions:
@@ 1. Do existing benchmarks really capture biases in the real world?
2. Can existing debias methods handle biases in the real world?
@@
To answer the second question, we propose two new biased distributions that better mirror biases in the real world. Together with biased distribution in existing benchmarks, we form a fine-grained evaluation framework for real-world debiasing. We focus on biased-auxiliary-model-based debias methods [12, 9, 15, 17, 18, 20], which is the predominant line of work in debiasing, and evaluate the methods with our evaluation framework. Empirical results show that existing methods fail to handle real-world biases. We further conducted an in-depth analysis of existing biased-auxiliary-model-based methods and found that the effectiveness of existing methods is reliant on the high bias prevalence of existing benchmarks and thus fails to handle real-world datasets with low bias prevalence.
Finally, according to our analysis on the biased distributions and existing debias methods [12, 9, 20], we introduce a simple yet effective enhancement to existing biased-auxiliary-model-based methods. Experiments show that our approach improves existing methods on both high and low bias prevalence datasets. To sum up, this work makes the following contributions:
  • Fine-grained analysis and evaluation framework. We propose a fine-grained framework for analyzing bias in datasets. Based on the framework we further introduce a systematic evaluation framework for real-world debiasing.
  • Theoretical insight. We derived the hidden strong assumptions held by existing benchmarks, and testify our observation that real-world biases are low in bias prevalence.
  • Principled approach. A principled approach is proposed based on our analysis, which can be easily applied to existing biased-auxiliary-model-based methods.
  • Empirical validation. Empirical results on multiple datasets and distributions not only validate our analysis but also demonstrate the superiority of the proposed approach.
@@ 2.1 Revisiting spurious correlation in datasets
In the area of spurious correlation debiasing, multiple synthetic [19, 12, 16] and semi-synthetic datasets [9, 12, 17] have been adopted to benchmark the effectiveness of the debias methods. Generally, those synthetic datasets first select a target attribute as the learning objective [16], e.g. object, and another spurious attribute that could potentially cause the learned model to be biased, e.g. background. Then, certain sub-groups jointly defined by the target and spurious attributes, e.g. water birds with water background, are emphasized, i.e. synthesized or sampled from real-world datasets with much higher probability (usually above 95%) than the others in the biased dataset construction process, causing the corresponding spurious feature and target feature to be spuriously correlated, e.g. water background correlated with water bird [16]. Specifically, one such dominating subgroup is selected for every possible value of the spurious attribute, forming a "diagonal distribution", as shown in Figure 1(a) and 1(c). However, it is critical to examine whether this pattern truly aligns with the complexities of real-world biases.
@@
We further investigated biases from the real world. COCO [21] dataset is a large-scale dataset collected from the internet and widely used in various vision tasks. COCO has been found to contain gender bias in web corpora [22], one of which is the spurious correlation between males and skateboards. The joint distribution of gender and Skateboard in COCO is plotted in Figure 1(b). COMPAS [23] dataset consists of the results of a commercial algorithm called COMPAS, used to assess a convicted criminal’s likelihood of reoffending. COMPAS dataset is widely known for its bias against African Americans and is widely used in the research of machine learning fairness [24, 25, 26, 27]. The joint distribution of Race and Risk Level in the COMPAS dataset is plotted in Figure 1(d). Note that although COMPAS is a tabular dataset, it genuinely reflects the biased distribution in the real world. It is quite obvious that the distribution of biases in existing benchmarks and real-world datasets diverges. More visualizations are shown in Appendix A. In the following subsection, we will further discuss how to measure their differences.
@@ 2.2 Previous measures of spurious correlation
Previous works [28, 19] measure spurious correlation according to the probability of a biased feature $a_{s}$ within the correlated class $a_{t}$: $Corr_{tcp}=P(y^{s}=a_{s}|y^{t}=a_{t})$. A higher value of the measure indicates a strong correlation.
@@
Some [22, 9, 29, 30] measure spurious correlation according to the probability of the correlated class $a_{s}$ within samples with biased feature $a_{t}$: $Corr_{scp}=P(y^{t}=a_{t}|y^{s}=a_{s})$. A higher value of the measure indicates a strong correlation.
@@ 2.3 The proposed analysis framework
Figure 2: Analysis of datasets bias in existing benchmarks and real-world datasets with the proposed framework. The bias magnitude and prevalence of real-world datasets are significantly smaller than that of existing benchmarks.
@@ 2.4 Observation on real-world biases
The magnitude of biases in real-world datasets is low. As shown in Figure 2(a), the magnitude of biases in real-world datasets is significantly lower than that in existing benchmarks. It is surprising to see how low the magnitude of biases in the COMPAS dataset is, yet still captured by models [26].
@@ 4 Methodology
In this section, we dive into how real-world biases would raise a challenge to existing debias methods. Specifically, we focus on debias with biased-auxiliary-model-based (DBAM) methods, which is the mainstream of work in the area of debiasing, and set up the problem in section 4.1. Then we claim the insufficiency of existing DBAM methods facing real-world biases in section 4.2. Finally, aiming at the insufficiency, we proposed our approach in section 4.3.
4.1 Problem setup
Debiasing with biased auxiliary model.
In recent years, DBAM methods have been widely studied [12, 9, 15, 17, 18, 20]. Based on the assumption that the spurious attributes are easier and learned more preferentially than the target attribute, these methods use a variant version of cross entropy(CE) that emphasizes easier samples, i.e. generalized cross entropy(GCE) [33], to train a biased auxiliary model $M_{b}$. Utilizing the bias captured by the biased auxiliary model, loss-based sample re-weighing scheme $W(x)$ [12] is applied to train the debiased model $M_{d}$:
@@ 4.1 Problem setup
where $(x,y)$ are samples from the training data and $CE(\cdot,\cdot)$ is the cross entropy loss.
4.2 Reliance of DBAM methods on high bias prevalence
We claim that the bias capture module of existing DBAM methods relies on the high bias magnitude of existing benchmarks, which causes failure in real-world debiasing. It is assumed by existing DBAM methods that the biased model $M_{b}$ predicts according to the bias within the training data, giving high loss to BC samples and low loss to BA samples [12, 9, 15, 17, 18, 20]. Existing works attribute this loss difference to the fact that spurious attributes are easier [12, 17], i.e. learned more preferentially by models, making BA the easy sample. While such a claim is true, we claim that the dominance of BA samples in the HP datasets is another vital causing factor of the loss difference, for dominant/major samples are learned more frequently than others, as shown in Figure 3(a).
However, on LP datasets, while BA samples are still easier to learn due to the biased feature, the dominant/major samples in the training data are no longer BA samples, but rather BN samples. This not only results in the loss difference between BA and BC samples decreasing but also causes low loss on BN samples, as shown in Figure 3(a). According to sample weighing scheme 4, such low loss on BN samples further leads to low weights for BN samples when training the debiased model, which is unintended because BN samples carry an abundant amount of knowledge concerning the target attribute without the interference of the spurious features. The overlooking of BN samples results in server utility degradation when existing DBAM methods are applied to LP datasets. We further empirically testify our claim in section 5.
@@ 4.3 Bias capture with feature destruction
To prune the side branch learning of the target features, it is intuitive to destroy the target feature and make them unlearnable when training the debiased model, as shown in Figure 3(b). Such action is practical because while the spurious features depend on the distribution of the training data, thus unknown, the target features we intend to learn are usually clear. Specifically, we can achieve this by applying target feature destructive data transformation when training the biased model:
$Loss_{b}=GCE(M_{b}(T_{fd}(x)),y)$
where $T_{fd}(\cdot)$ is the feature destruction transformation. As an example, in visual recognition tasks, the shape of objects is a basic element of human visual perception [34]. Therefore, the destruction of shape when capturing bias from visual recognition datasets is a feasible approach. Specifically, various approaches can be applied to destroy the shape feature in data [35], e.g. pixel-mixed and patch-mixed image transformation. Here we adopt pixel-mixed image transformation following [35].
@@ 5 Experiments
In this section, based on our analysis in section 2, we first propose our fine-grained evaluation framework for real-world debiasing. Then, we design our experiments to answer the following questions: 1) How do existing DBAM methods and DiD perform on real-world biases? 2) Does DiD really emphasize BN samples as we intend? 3) How do debias methods perform on unbiased datasets? 4) How sensitive is DiD to the hyper-parameters? 5) How do the magnitude and prevalence of bias in datasets affect debiasing?
5.1 Fine-grained evaluation framework
As discussed in section 2, biases present in current synthetic benchmarks do not accurately reflect those in the real world. To mirror real-world biases more closely, we introduce two new bias distributions. Together with high magnitude high prevalence (HMHP) biased distribution in existing benchmarks [12, 16, 9, 17], we form our fine-grained evaluation framework for real-world debiasing.
Low Magnitude Low Prevalence (LMLP) Bias.
Inspired by the distribution of the COMPAS [23] dataset shown in Figure 2(a), bias in the real world might be low in both magnitude and prevalence.
High Magnitude Low Prevalence (HMLP) Bias.
Note that datasets with low magnitude high prevalence (LMHP) bias do not exist due to the fact that high bias magnitude is the premise of high bias prevalence. Also, the bias distributions proposed are applicable to existing synthetic and semi-synthetic datasets mentioned in section 2. More details on how biased distributions are defined and how to synthesize datasets with given distributions can be found in Appendix B. In this section, we utilize the basic setting of Colored MNIST [19] and Corrupted CIFAR10 [12] for our experiments, details of which can be found in Appendix D.
@@ 5.2 Experimental settings
Following previous works [12, 9, 20], we adopt the accuracy of BC samples (BC) and the average accuracy on the balanced test set (Avg) as the evaluation metrics. We adopt 5 baselines, covering classic and recently proposed DBAM methods. ERM directly applies standard training on the biased datasets. LfF [12] is the first work to propose the DBAM framework. DisEnt [9] disentangles bias and intrinsic features and applies feature augmentation when training the debiased model. LfF BE and DisEnt BE were recently proposed by Lee et al. [20], and is based on LfF and DisEnt, respectively. These two methods strengthen the bias capture process by applying an ensemble of biased classifiers. Detailed settings can be found in Appendix D.
@@ 5.3 Main results
As shown in Table 1, while performing decently on HMHP distributed datasets, existing DBAM methods [12, 9, 20] fail to handle both LMLP and HMLP biases. Generally, for existing DBAM methods, the accuracy of BC samples and the average accuracy on the balanced test set is lower than ERM baseline. This indicates the failure of exiting DBAM methods on the task of debiasing with low bias prevalence datasets, sacrificing utility, i.e. average accuracy, without improving worst group performance, i.e. BC accuracy. Detailed results of existing DBAM methods including the accuracy of all types of samples can be found in Appendix E. We further tested the effectiveness of DiD by combining DiD with existing DBAM methods [12, 9, 20]. As shown in Table 1, after combining DiD, the BC sample accuracy and average accuracy on a balanced test set both improve on existing HMHP benchmarks. On LMLP and HMLP datasets, the superiority of DiD is even more prominent, where BC and average accuracy both improve significantly.
Table 1: The performance of our approach is presented in absolute accuracy increase of existing methods. Results show that existing DBAM methods perform poorly on LP distributions, yet our method effectively boosts the performance of existing DBAM methods.
Colored MNIST Corrupted CIFAR10 LMLP HMLP HMHP LMLP HMLP HMHP Algorithm BC Avg BC Avg BC Avg BC Avg BC Avg BC Avg ERM 91.1 91.7 85.2 89.8 48.5 53.4 62.5 64.3 55.9 65.1 29.4 35.4 LfF 68.4 69.7 58.0 63.3 65.6 64.6 55.0 55.4 47.7 54.1 35.3 39.0 + Ours +22.6 +21.4 +32.6 +25.8 +1.3 +3.4 +7.0 +7.3 +7.1 +8.9 +1.8 +2.5 LfF BE 83.6 83.5 80.0 82.3 66.9 67.6 52.1 54.0 51.0 54.0 31.5 36.6 + Ours +5.7 +6.1 +9.1 +4.9 -0.5 +0.7 +1.1 +0.2 -0.8 +0.1 +1.4 +0.8 DisEnt 73.9 74.9 66.5 72.2 68.3 67.4 55.5 56.1 52.5 54.5 36.0 39.5 + Ours +17.2 +16.5 +22.0 +16.8 +0.8 +3.1 +5.4 +5.9 +2.8 +7.1 +3.0 +3.3 DisEnt BE 81.1 81.0 77.6 80.2 67.5 68.5 56.6 57.2 49.1 56.3 34.2 38.6 + Ours +8.7 +9.0 +11.7 +5.5 +2.0 +2.5 +4.3 +4.2 +4.9 +5.1 +3.5 +3.2

### v3

# Rethinking Debiasing: Real-World Bias Analysis and Mitigation

@@ Abstract
Spurious correlations in training data significantly hinder the generalization capability of machine learning models when faced with distribution shifts in real-world scenarios. To tackle the problem, numerous debiasing approaches have been proposed and benchmarked on datasets intentionally designed with severe biases. However, it remains to be asked: 1. Do existing benchmarks really capture biases in the real world? 2. Can existing debiasing methods handle biases in the real world? To answer the questions, we revisit biased distributions in existing benchmarks and real-world datasets, and propose a fine-grained framework for analyzing dataset bias by disentangling it into the magnitude and prevalence of bias. We empirically and theoretically identify key characteristics of real-world biases poorly represented by existing benchmarks. We further introduce two novel biased distributions to bridge this gap, forming a systematic evaluation framework for real-world debiasing. With the evaluation framework, we focus on the practical setting of debiasing w/o bias supervision and find existing methods incapable of handling real-world biases. Through in-depth analysis, we propose a simple yet effective approach that can be easily applied to existing debiasing methods, named Debias in Destruction (DiD). Empirical results on real-world datasets in both image and language modalities demonstrate the superiority of DiD, improving the performance of existing methods on all types of biases.
@@ 1 Introduction
To tackle the problem, various debiasing methods [13, 12, 14, 9, 15, 16, 17, 18] have been proposed in recent years. And the effectiveness of the methods is benchmarked with synthetic [19, 12, 16] and semi-synthetic [9, 12, 17] (referred as "real-world dataset" in previous works) datasets designed to be severely biased. However, while these benchmarks are indeed biased, they are rough and lack thorough consideration of how data is truly biased in the real world. This raises two questions:
@@ 1. Do existing benchmarks really capture biases in the real world?
2. Can existing debiasing methods handle biases in the real world?
@@
As for the second question, due to the complexity of real-world biases, debiasing methods should be capable of handling various types of biased distributions rather than solely on the narrowed distribution in existing benchmarks. Thus, we introduce two new biased distributions inspired by real-world applications as a complement to the biased distribution in existing benchmarks, forming a systematic evaluation framework for real-world debiasing. We focus on debiasing methods w/o bias supervision[12, 9, 16, 15, 17, 18, 20], which is more practical as bias feature is expensive to annotate and sometimes even hard to notice [21]. Those methods generally involves a biased auxiliary model to capture the bias, along with techniques to learn a debiased model with the captured bias. (See the related work section in Appendix F for details) We condense them into a paradigm called debiasing with biased auxiliary model (DBAM). Our empirical results show that existing methods fail to handle real-world biases under various biased distributions.
We further conducted an in-depth analysis of existing methods and found that their effectiveness is reliant on the high bias prevalence of existing benchmarks and thus fails to handle real-world datasets with low bias prevalence. Finally, based on our analysis, we introduce a simple yet effective enhancement to existing methods. Experiments on real-world datasets in both image and language modalities show that our approach significantly boosts the capability of existing methods in handling real-world biases, improving their performances on both high and low bias prevalence datasets. To sum up, this work makes the following contributions:
  • Identify key characteristics of real-world biases. We propose a fine-grained framework for analyzing bias in datasets. Based on the framework, we empirically and theoretically identified two key characteristics of real-world biases, previously overlooked. We then propose a systematic evaluation framework inspired by real-world applications for real-world debiasing.
  • Uncover critical failure mode of existing unsupervised debiasing methods. Evaluation of existing methods shows that they fail at handling real-world biases. We further prove that the failure is due to their reliance on unrealistic assumptions on the data distribution.
  • Principled approach. A simple yet effective approach is proposed based on our analysis, which can be easily applied to existing unsupervised debiasing methods.
  • Empirical validation. Extensive empirical results on multiple real-world datasets, distributions, and modalities not only validate our analysis but also demonstrate the superiority of the proposed approach.
@@ 2.1 Revisiting spurious correlation in datasets
In the area of spurious correlation debiasing, multiple synthetic [19, 12, 16] and semi-synthetic datasets [9, 12, 17] have been adopted to benchmark the effectiveness of the debiasing methods. Generally, those synthetic datasets first select a target attribute as the learning objective [16], e.g. object, and another spurious attribute that could potentially cause the learned model to be biased, e.g. background. Then, certain sub-groups jointly defined by the target and spurious attributes, e.g. water birds with water background, are emphasized, i.e. synthesized or sampled from real-world datasets with much higher probability (usually above 95%) than the others in the biased dataset construction process, causing the corresponding spurious feature and target feature to be spuriously correlated, e.g. water background correlated with water bird [16]. Specifically, one such dominating subgroup is selected for every possible value of the spurious attribute, forming a "diagonal distribution", as shown in Figure 1(a) and 1(c). However, it is critical to examine whether this pattern truly aligns with the complexities of real-world biases.
@@
We further investigated biases from the real world. COCO [22] dataset is a large-scale dataset collected from the internet and widely used in various vision tasks. COCO has been found to contain gender bias in web corpora [23], one of which is the spurious correlation between males and skateboards. The joint distribution of gender and Skateboard in COCO is plotted in Figure 1(b). COMPAS [24] dataset consists of the results of a commercial algorithm called COMPAS, used to assess a convicted criminal’s likelihood of reoffending. COMPAS dataset is widely known for its bias against African Americans and is widely used in the research of machine learning fairness [25, 26, 27, 28]. The joint distribution of Race and Risk Level in the COMPAS dataset is plotted in Figure 1(d). Note that although COMPAS is a tabular dataset, it genuinely reflects the biased distribution in the real world. It is quite obvious that the distribution of biases in existing benchmarks and real-world datasets diverges. Additonally, CelebA [29] is another real-world image dataset. CivilComments-WILDS (CCW) [3] and MultiNLI [30] are also real-world datasets in the NLP domain. More visualizations of these additional datasets are shown in Appendix A. In the following subsection, we will further discuss how to measure their differences.
@@ 2.2 Previous measures of spurious correlation
Previous works [31, 19] measure spurious correlation according to the probability of a biased feature $a_{s}$ within the correlated class $a_{t}$: $Corr_{tcp}=P(y^{s}=a_{s}|y^{t}=a_{t})$. A higher value indicates a strong correlation.
@@
Some [23, 9, 32, 33] measure spurious correlation according to the probability of the correlated class $a_{t}$ within samples with biased feature $a_{s}$: $Corr_{scp}=P(y^{t}=a_{t}|y^{s}=a_{s})$. A higher value of the measure indicates a strong correlation.
@@ 2.3 The proposed analysis framework
Figure 2: With our analysis framework, we can see that the bias magnitude and prevalence of real-world datasets are significantly smaller than that of existing benchmarks.
@@ 2.4 Observation on real-world biases
The magnitude of biases in real-world datasets is low. As shown in Figure 2(a), the magnitude of biases in real-world datasets is significantly lower than that in existing benchmarks, consistent across various modalities. It is surprising to see how low the magnitude of biases in the real-world dataset is, yet still captured by models [27].
@@
2.5 Systematic evaluation framework for real-world debiasing
Based on our analysis, we further introduce two novel bias distributions inspired from real-world applications. Together with high magnitude high prevalence (HMHP) distribution in existing benchmarks [12, 16, 9, 17], we form a systematic evaluation framework for real-world debiasing.
Low Magnitude Low Prevalence (LMLP) Bias.
Inspired by the distribution of the COMPAS [24] dataset shown in Figure 2(a), bias in the real world might be low in both magnitude and prevalence. To take it even further, we should not even assume the dataset is biased at all when applying debiasing methods, because we usually lack such information in practice. Thus, unbiased data distribution can be considered as a special case of the distribution.
High Magnitude Low Prevalence (HMLP) Bias.
Note that datasets with low magnitude high prevalence (LMHP) bias do not exist due to the fact that high bias magnitude is the premise of high bias prevalence. Also, the bias distributions proposed are applicable to existing synthetic and semi-synthetic datasets mentioned in section 2. More details on how biased distributions are defined and how to synthesize datasets with given distributions can be found in Appendix B. In section 5, We further examine existing methods and our proposed method under real-world biases with the evaluation framework.
@@ 4 Methodology
In this section, we dive into how real-world biases would raise a challenge to existing debiasing methods. Specifically, we first summarize and condense existing unsupervised debiasing methods into the debiasing with biased auxiliary model (DBAM) paradigm in section 4.1. Then we claim the insufficiency of existing DBAM paradigm when facing real-world biases in section 4.2. Finally, aiming at the insufficiency, we proposed our approach in section 4.3.
4.1 Debiasing with biased auxiliary model
In recent years, research in the field of debiasing has been more focused on the practical setting of debiasing w/o bias supervision. Though different in technical details, they generally adopt a biased auxiliary model to capture the bias, followed by techniques to learn a debiased model with the captured bias. The bias capture process is based on the assumption that the spurious attributes are easier and learned more preferentially than the target attribute, thus bias could be captured by an auxiliary model $M_{b}$ w/o bias supervision. To utilize $M_{b}$ for debiasing, sample re-weighing schemes have been widely incorporated into the training of the debiased model $M_{d}$, one common implementation of which is loss-based re-weighing $W(x)$ [12]:
@@ 4.1 Debiasing with biased auxiliary model
where $(x,y)$ are samples from the training data and $CE(\cdot,\cdot)$ is the cross entropy loss. We note that the above design is introduced because it has been widely adopted by various algorithms, making it a solid entry point for our subsequent analysis. DBAM paradigm is our condensation of existing unsupervised debiasing methods, and we do not limit its scope with the above re-weighing design $W(x)$.
4.2 DBAM paradigm relies on high bias prevalence
We claim that the bias capture module of the DBAM paradigm relies on the high bias magnitude of existing benchmarks, which causes failure in real-world debiasing. It is assumed by the DBAM paradigm that the biased model $M_{b}$ predicts according to the bias within the training data, giving high loss to BC samples and low loss to BA samples [12, 9, 15, 17, 18, 20]. Existing works attribute this loss difference to the fact that spurious attributes are easier [12, 17], i.e., learn more preferentially by models, making BA the easy sample. While such a claim is true, we claim that the dominance of BA samples in the HP datasets is another vital causing factor of the loss difference, for dominant/major samples are learned more frequently than others, as shown in Figure 3(a).
However, on LP datasets, while BA samples are still easier to learn due to the biased feature, the dominant/major samples in the training data are no longer BA samples, but rather BN samples. This not only results in the loss difference between BA and BC samples decreasing but also causes low loss on BN samples, as shown in Figure 3(a). According to sample weighing scheme 4, such low loss on BN samples further leads to low weights for BN samples when training the debiased model, which is unintended as BN samples carry an abundant amount of knowledge concerning the target attribute without the interference of the spurious features. Overlooking BN samples results in server utility degradation when DBAM methods are applied to LP datasets. We empirically testify our claim in section 5.
@@ 4.3 Bias capture with feature destruction
To prune the side branch learning of the target features, it is intuitive to destroy the target feature and make them unlearnable when training the biased model, as shown in Figure 3(b). Such action is practical because the target features we intend to learn are usually clear, and no information on the biased feature is required. Specifically, we can achieve this by applying target feature destructive data transformation when training the biased model:
$Loss_{b}^{DiD}=Loss_{b}(M_{b}(T_{fd}(x)),y)$
where $Loss_{b}$ is the original loss for training the biased model (e.g. $CE$), and $T_{fd}(\cdot)$ is the feature destruction transformation. As an example, in visual recognition tasks, the shape of objects is a basic element of human visual perception [36]. Therefore, the patch-shuffle destruction of shape [37] when capturing bias from visual recognition datasets is a feasible approach. As for NLP tasks, we adopt a word shuffling approach which we will elaborate on in Appendix D.5.
@@ 5 Experiments
In this section, with the systematic evaluation framework proposed in section 2.5, we design our experiments to answer the following questions: 1) How do existing DBAM methods and DiD perform on real-world biases? 2) Does DiD really emphasize BN samples as we intend? 3) How do debiasing methods perform on unbiased datasets? 4) How do the magnitude and prevalence of bias in datasets affect debiasing? 5) How sensitive is DiD to the hyper-parameters? (see Appendix E.2) 6) Is DiD effective on bias detection tasks as well? (see Appendix E.3) More analytical results in Appendix E.
@@ 5.1 Experimental settings
Metrics.
Following previous works [12, 9, 20], we adopt the accuracy of BC samples (BC), the average accuracy on the balanced test set (Avg), and the worst group accuracy (Worst Acc.) as the evaluation metrics. We note that DiD is not a self-contained method but rather a plug-in module for existing debiasing methods to improve from their original performances. Thus, the effectiveness of DiD should be measured as the performance gain from the base method for debiasing.
Datasets.
We adopt 8 datasets in various modalities for evaluation. Specifically, we adopt the basic setting of Colored MNIST [19] and Corrupted CIFAR10 [12] to implement the distributions within the proposed systematic evaluation framework. We also evaluated our method on more existing synthetic benchmarks who is more complex in terms of the target and spurious feature: BAR [12], NICO [15], and WaterBirds [38]. We also adopt 2 real-world NLP datasets MultiNLI [30] and CivilComments-WILDS [3], and 1 real-world image dataset CelebA [29].
Baselines.
We adopt 7 baselines, covering classic and recently proposed DBAM methods. ERM directly applies standard training on the biased datasets. LfF [12] is the first work in the DBAM paradigm. DisEnt [9] disentangles bias and intrinsic features and applies feature augmentation when training the debiased model. BE LfF and BE DisEnt were recently proposed by Lee et al. [20], and is based on LfF and DisEnt, respectively. JTT [16] is a classic DBAM method adapted to both the image and NLP domain. Group DRO [38] is a classic debiasing method with bias supervision used as an upper bound. Detailed settings can be found in Appendix D.
@@ 5.2 Main results
Evaluation on various bias distributions.
As shown in Table 1, while performing decently on HMHP distributed datasets, existing methods [12, 9, 20] fail to handle both LMLP and HMLP biases. Generally, for existing methods, the accuracy of BC samples and the average accuracy on the balanced test set is lower than the ERM baseline. This indicates the failure of exiting DBAM paradigm on the task of debiasing with low bias prevalence datasets, sacrificing utility, i.e. average accuracy, without improving worst group performance, i.e. BC accuracy. We further tested the effectiveness of DiD by combining DiD with existing DBAM methods [12, 9, 20]. As shown in Table 1, when combined with DiD, the BC and average accuracy both improve on existing HMHP benchmarks. On LMLP and HMLP datasets, the superiority of DiD is even more prominent, where BC and average accuracy both improve significantly, achieving an average of +16.6 and +11.7 for LfF and DisEnt respectively.
Table 1: The performance of our approach is presented in absolute accuracy increase of existing methods. Results show that existing DBAM methods perform poorly on LP distributions, yet our method effectively boosts the performance of existing methods across all types of biases.
Colored MNIST Corrupted CIFAR10 LMLP HMLP HMHP LMLP HMLP HMHP Algorithm BC Avg BC Avg BC Avg BC Avg BC Avg BC Avg ERM 91.1 91.7 85.2 89.8 48.5 53.4 62.5 64.3 55.9 65.1 29.4 35.4 LfF 68.4 69.7 58.0 63.3 65.6 64.6 55.0 55.4 47.7 54.1 35.3 39.0 + Ours +22.6 +21.4 +32.6 +25.8 +1.3 +3.4 +7.0 +7.3 +7.1 +8.9 +1.8 +2.5 DisEnt 73.9 74.9 66.5 72.2 68.3 67.4 55.5 56.1 52.5 54.5 36.0 39.5 + Ours +17.2 +16.5 +22.0 +16.8 +0.8 +3.1 +5.4 +5.9 +2.8 +7.1 +3.0 +3.3 BE LfF 83.6 83.5 80.0 82.3 66.9 67.6 52.1 54.0 51.0 54.0 31.5 36.6 + Ours +5.7 +6.1 +9.1 +4.9 -0.5 +0.7 +1.1 +0.2 -0.8 +0.1 +1.4 +0.8 BE DisEnt 81.1 81.0 77.6 80.2 67.5 68.5 56.6 57.2 49.1 56.3 34.2 38.6 + Ours +8.7 +9.0 +11.7 +5.5 +2.0 +2.5 +4.3 +4.2 +4.9 +5.1 +3.5 +3.2
Table 2: Our approach consistently demonstrated the effectiveness on real-world datasets in both image and language modality. Group DRO is a supervised debiasing method, acting as an upper bound for worst-group accuracy.
Bias supervision? MultiNLI CivilComments-WILDS CelebA Avg Worst Acc. Avg Worst Acc. Avg Worst Acc. ERM No 80.1 76.41 92.06 50.87 95.75 45.56 JTT No 80.51 73.02 91.25 59.49 80.49 73.13 +Ours No +1.06 +2.71 +0.38 +6.41 +6.43 +8.50 Group DRO Yes 82.11 78.67 83.92 80.20 91.96 91.49
Evaluation on complex visual features.
As shown in Table 4, our approach is not merely effective under the setting of Colored MNIST and Corrupted CIFAR10, but rather consistently demonstrating its superiority on datasets with more complex sets of target and spurious features. This shows the adaptability of DiD to more sophisticated visual data. Refer to Appendix D.2 for the metrics used.
Evaluation on real-world datasets in various modalities.
We choose JTT as the baseline for this part of the experiment for it is a classic method adopted to both the image and NLP domain. As shown in Table 2, our approach is consistently effective on real-world datasets in various modalities, further demonstrating its generalizability.